# D-52056 Aachen, Germany

### v2

@@ 6 Differential fiducial cross-section distributions
The observables discussed so far share a common characteristic, namely their NLO QCD corrections are rather stable and of the order of $20\%$. Moreover, the corresponding NLO uncertainties are of a similar magnitude. Finally, the NLO uncertainty bands consistently coincide with the corresponding LO ones. However, there are certain types of observables for which NLO QCD corrections can become very large, particularly, in some phase-space regions and the NLO predictions might no longer fall within the LO uncertainty bands. For illustrative purposes, in Figure 8, we show differential cross-section distributions for the transverse momentum of the $b_{1}b_{2}$ and $b_{1}b_{2}b_{3}b_{4}$ systems. At LO, in the rest frame of the (anti-)top quark the maximum transverse momentum of the resulting $b$-jet is given by $p_{T,\,b}=(m_{t}^{2}-m_{W}^{2})/2m_{t}\approx 67.5$ GeV. In order to produce $b$-jets with $p_{T,\,b}>67.5$ GeV they need to be boosted via the $p_{T}$ of the parent particle. Consequently, at LO the $b\bar{b}b\bar{b}$ system with high $p_{T}$ is kinematically strongly disfavoured. At NLO, on the other hand, the $t\bar{t}t\bar{t}$ system can acquire large transverse momentum by recoiling against the extra light jet, resulting in a significant enhancement of events at large values of $p_{T,\,b_{1}b_{2}b_{3}b_{4}}$. Indeed, we can observe that the differential ${\cal K}$-factor starts to rapidly increase for $p_{T,\,b_{1}b_{2}b_{3}b_{4}}\gtrsim 270$ GeV and NLO QCD predictions can be as much as $2.4$ times larger than the LO ones in the tails of the $p_{T,\,b_{1}b_{2}b_{3}b_{4}}$ distribution. We can also notice that the NLO theoretical uncertainties in those phase-space regions substantially increase and are of the order of $50\%$. Finally, the NLO results are not within the corresponding LO uncertainties that are at the level of $76\%$. Similar but less pronounced effects can also be observed for the $p_{T,\,b_{1}b_{2}}$ observable in the same plotted range. In this case, the NLO QCD corrections increase maximally to $45\%$, while NLO uncertainties to $30\%$. To illustrate the strong dependence of these two observables on extra jet radiation, we can analyse them again, this time also applying a veto on the additional light- or $b$-jet. The difference between the NLO cross section without and with this jet veto is given by a LO calculation, since in the latter case the existence of the five well-separated jets is demanded. While the four $b$-jets still need to have $p_{T,\,b}>25$ GeV, the additional jet, if resolved and passing all the cuts that are also required for the $b$-jet, should additionally have $p_{T,\,j}<p_{T}^{veto}$, with $p_{T}^{veto}=(25,50,100,200,300,500)$ GeV, where $j=g,u,d,c,s,b$. Jet vetos significantly affect the normalisation of differential cross-section distributions, consequently, we study only their normalised versions. In Figure 9 we show the normalised LO and NLO QCD differential cross-section distributions for $p_{T,\,b_{1}b_{2}b_{3}b_{4}}$ and $p_{T,\,b_{1}b_{2}}$. Also depicted are the NLO QCD results with the jet veto of $p_{T}^{veto}=(25,50,100,200,300,500)$ GeV. For various values of the $p_{T}^{veto}$ cut large differences in the differential factor ${\cal K}$ can be observed. The substantial higher-order QCD corrections that we could previously see in the tails of the two distributions have now been reduced to just a few per cent. For $p_{T}^{veto}=25$ GeV we obtain negative values of the NLO cross section for $p_{T}\geq 470$ GeV for both observables. Thus, this particular value is too low and should not be considered in experimental analyses. It should be noted, however, that the jet veto might substantially increase the size of the NLO uncertainty bands. For example, for $p_{T}^{veto}=50$ GeV the NLO uncertainties increase by a factor of $3-7$ depending on the observable. The observed large scale dependence for the more exclusive NLO results for $p_{T,\,b_{1}b_{2}}$ and $p_{T,\,b_{1}b_{2}b_{3}b_{4}}$ gives cause for concern, as jet vetoes are widely used in experimental analyses at the LHC. For example, vetoing the additional jet activity is crucial for suppressing various SM background processes or in various searches for new physics effects. Indeed, in many analyses carried out by the ATLAS and CMS collaborations the most common jet veto scheme comprises imposing a maximum transverse momentum cut on anti-$k_{T}$ jets. Therefore, in such cases, in order to reduce theoretical uncertainties and to improve the convergence of perturbative predictions, approaches similar to those described in e.g. Ref. Berger:2010xi ; Stewart:2010tn ; Becher:2012qa ; Banfi:2012yh ; Banfi:2012jm ; Michel:2018hui ; Campbell:2023cha would have to be used.
@@
In the end, we examine the previously defined three approaches $d\sigma^{\rm NLO}_{\rm exp}/dX$, $d\sigma^{\rm NLO}_{\rm full}/dX$ and $d\sigma^{\rm NLO}_{\rm LO_{dec}}/dX$ as a function of the observable $X$, where $X=M_{b_{1}b_{2}},\,p_{T,\,\ell_{1}},\,p_{T,\,b_{1}}$ and $p_{T,\,b_{1}b_{2}}$. These results are shown in Figure 11. The upper panels show the absolute NLO QCD predictions for the three approaches together with their corresponding theoretical uncertainties as obtained from the scale dependence. The two middle panels display the ratios of $d\sigma^{\rm NLO}_{\rm full}/dX$ and $d\sigma^{\rm NLO}_{\rm LO_{dec}}/dX$ to the $d\sigma^{\rm NLO}_{\rm exp}/dX$ result, together with their relative uncertainties. The lowest panels compare again the relative size of NLO scale uncertainties for the three approaches, this time, however, the findings are normalised to their corresponding NLO results for $\mu_{0}=E_{T}/4$. For the three observables studied $M_{b_{1}b_{2}},\,p_{T,\,\ell_{1}}$ and $p_{T,\,b_{1}}$, our observations and conclusions are very similar. The differences between the full results and the expanded ones are of the order of $10\%$, which should be compared to the corresponding NLO uncertainties that are in the range of $15\%-25\%$. On the other hand, NLO QCD corrections in top-quark decays are consistently at the level of $10\%$. Larger effects could be observed for $p_{T,\,b_{1}b_{2}}$, especially in the tail of the distribution. In this case the differences between $d\sigma^{\rm NLO}_{\rm exp}/dp_{T,\,b_{1}b_{2}}$ and $d\sigma^{\rm NLO}_{\rm full}/dp_{T,\,b_{1}b_{2}}$ increase to $20\%$. However, the NLO uncertainties are also increasing and are now in the range of $25\%-35\%$. Nevertheless, the size of higher-order QCD effects in top-quark decays remains similar to the case of the other three observables. In short, the overall picture is rather similar to that already observed at the integrated fiducial cross-section level. Firstly, the magnitude of the NLO QCD uncertainties increases when higher-order QCD corrections to top-quark decays are neglected. The latter effects are consistently at the $10\%$ level. Secondly, the differences between the expanded and unexpanded NLO results are within the obtained NLO uncertainties. However, the lack of the expansion of the $1/\Gamma_{t}$ factor in $d\sigma^{\rm NLO}_{\rm full}/dX$ significantly reduces the size of the NLO uncertainties for all the observables that we have examined.
7 Summary
@@
In the next step, we have compared the following NLO predictions: $d\sigma^{\rm NLO}_{\rm exp}/dX$, $d\sigma^{\rm NLO}_{\rm full}/dX$ and $d\sigma^{\rm NLO}_{\rm LO_{dec}}/dX$ as a function of various observables $X$. By analysing these predictions we have concluded that the overall pattern is rather similar to that which has already been observed at the integrated fiducial cross-section level. Namely, the magnitude of the NLO uncertainties increases when NLO QCD corrections to top-quark decays are neglected. The higher-order corrections in top-quark decays are rather stable and of the order of $10\%$. In addition, the difference between the expanded and unexpanded NLO results is within the obtained NLO scale uncertainties. Nonetheless, the lack of the expansion of the $1/\Gamma_{t}$ term in $d\sigma^{\rm NLO}_{\rm full}/dX$, significantly reduces the size of the NLO uncertainties for all the observables that we have scrutinised.
@@
In addition, it would be advantageous to examine the so-called Matrix Element Corrections (MEC) to the Pythia8 parton shower program Norrbin:2000uu that can upgrade the accuracy of top-quarks decays to approximately NLO QCD in the shower evolution. Quite recently, a study has been carried out taking into account the MEC for the simplest case of $pp\to t\bar{t}$ production in di-lepton top-quark decay channel Frixione:2023hwz . Another study has just been performed for the $pp\to t\bar{t}W^{\pm}$ process for the two same-sign leptons and jets $(2SS\ell)$ as well as three-lepton $(3\ell)$ final states Frederix:2024psm . Unfortunately, such analyses are not yet available for the $pp\to t\bar{t}t\bar{t}+X$ process. Although NLO plus parton shower predictions in the Powheg Box and Mc@normal-@@@Nlo framework for $pp\to t\bar{t}t\bar{t}+X$ production in the $\ell+jets$ channel are available in the literature Jezo:2021smh , the MEC to the top-quark decays are not included there. As this is a much more complicated process with two $t\bar{t}$ pairs and at least 12 final state particles, it would be interesting to test the validity of such approximate methods in such a complex environment. Another very important issue to look at when using the MEC is spin correlations. Having fixed-order predictions that describe top-quark decays and spin correlations with NLO QCD accuracy, we are in an excellent position to test both aspects. Again, we plan to perform such a comparison in the near future.
@@
The obtained NLO QCD results for the $pp\to t\bar{t}t\bar{t}+X$ process are especially important for the High-Luminosity phase of the LHC (HL-LHC). The upgrade is currently in progress and ATLAS and CMS experiments are expected to start taking data from the beginning of 2029. The HL-LHC will allow us to study $pp\to t\bar{t}t\bar{t}$ production in the $4\ell$ channel in greater detail. We point out that our conclusions regarding the magnitude of NLO QCD corrections, the significance of higher-order effects in top-quark decays or the size of NLO uncertainties will remain the same if the center-of-mass energy of the LHC is increased from the current value to $14$ TeV. Finally, we note that, despite its relatively small cross section, a good theoretical control over the $pp\to t\bar{t}t\bar{t}$ process in the $4\ell$ channel is phenomenologically very relevant. Indeed, the NLO QCD uncertainties that we have estimated for this channel are at the level of $15\%-20\%$. For comparison, the projected uncertainties for the HL-LHC, combining ATLAS and CMS, are in the range of $\delta=1/\sqrt{N}=15\%-18\%$. The latter uncertainties are estimated on the basis of an integrated luminosity of $(3-4)$ ${\rm ab}^{-1}$, which is planned for the HL-LHC ZurbanoFernandez:2020cco . Thus, at the HL-LHC, among other things, we will gain sensitivity to higher-order effects in top-quark decays in the $pp\to t\bar{t}t\bar{t}+X$ process. However, even now, pending new results from HL-LHC, the NLO QCD results obtained in this article are highly relevant for various preparatory studies. Proper modelling of differential cross-section distributions, for example, is an invaluable contribution to the many Machine Learning studies that are currently being conducted for this process. We also believe that the NLO QCD analysis for the SM $pp\to t\bar{t}t\bar{t}+X$ (background) process at the LHC is a necessary step towards a correct interpretation of possible signals of new physics that may arise in the $4\ell$ channel.

### v3

@@ 6 Differential fiducial cross-section distributions
The observables discussed so far share a common characteristic, namely their NLO QCD corrections are rather stable and of the order of $20\%$. Moreover, the corresponding NLO uncertainties are of a similar magnitude. Finally, the NLO uncertainty bands consistently coincide with the corresponding LO ones. However, there are certain types of observables for which NLO QCD corrections can become very large, particularly, in some phase-space regions and the NLO predictions might no longer fall within the LO uncertainty bands. For illustrative purposes, in Figure 8, we show differential cross-section distributions for the transverse momentum of the $b_{1}b_{2}$ and $b_{1}b_{2}b_{3}b_{4}$ systems. At LO, in the rest frame of the (anti-)top quark the maximum transverse momentum of the resulting $b$-jet is given by $p_{T,\,b}=(m_{t}^{2}-m_{W}^{2})/2m_{t}\approx 67.5$ GeV. In order to produce $b$-jets with $p_{T,\,b}>67.5$ GeV they need to be boosted via the $p_{T}$ of the parent particle. Consequently, at LO the $b\bar{b}b\bar{b}$ system with high $p_{T}$ is kinematically strongly disfavoured. At NLO, on the other hand, the $t\bar{t}t\bar{t}$ system can acquire large transverse momentum by recoiling against the extra light jet, resulting in a significant enhancement of events at large values of $p_{T,\,b_{1}b_{2}b_{3}b_{4}}$. Indeed, we can observe that the differential ${\cal K}$-factor starts to rapidly increase for $p_{T,\,b_{1}b_{2}b_{3}b_{4}}\gtrsim 270$ GeV and NLO QCD predictions can be as much as $2.4$ times larger than the LO ones in the tails of the $p_{T,\,b_{1}b_{2}b_{3}b_{4}}$ distribution. We can also notice that the NLO theoretical uncertainties in those phase-space regions substantially increase and are of the order of $50\%$. Finally, the NLO results are not within the corresponding LO uncertainties that are at the level of $76\%$. Similar but less pronounced effects can also be observed for the $p_{T,\,b_{1}b_{2}}$ observable in the same plotted range. In this case, the NLO QCD corrections increase maximally to $45\%$, while NLO uncertainties to $30\%$. To illustrate the strong dependence of these two observables on extra jet radiation, we can analyse them again, this time also applying a veto on the additional light- or $b$-jet. The difference between the NLO cross section without and with this jet veto is given by a LO calculation, since in the latter case the existence of the five well-separated jets is demanded. While the four $b$-jets still need to have $p_{T,\,b}>25$ GeV, the additional jet, if resolved and passing all the cuts that are also required for the $b$-jet, should additionally have $p_{T,\,j}<p_{T}^{veto}$, with $p_{T}^{veto}=(25,50,100,200,300,500)$ GeV, where $j=g,u,d,c,s,b$. Jet vetos significantly affect the normalisation of differential cross-section distributions, consequently, we study only their normalised versions. In Figure 9 we show the normalised LO and NLO QCD differential cross-section distributions for $p_{T,\,b_{1}b_{2}b_{3}b_{4}}$ and $p_{T,\,b_{1}b_{2}}$. Also depicted are the NLO QCD results with the jet veto of $p_{T}^{veto}=(25,50,100,200,300,500)$ GeV. For various values of the $p_{T}^{veto}$ cut large differences in the differential factor ${\cal K}$ can be observed. The substantial higher-order QCD corrections that we could previously see in the tails of the two distributions have now been reduced to just a few per cent. For $p_{T}^{veto}=25$ GeV we obtain negative values of the NLO cross section for $p_{T}\geq 470$ GeV for both observables. This pathological behaviour of fixed-order calculations can be cured by using shower effects or analytic resummation. It should be noted, however, that the jet veto might substantially increase the size of the NLO uncertainty bands. For example, for $p_{T}^{veto}=50$ GeV the NLO uncertainties increase by a factor of $3-7$ depending on the observable. The observed large scale dependence for the more exclusive NLO results for $p_{T,\,b_{1}b_{2}}$ and $p_{T,\,b_{1}b_{2}b_{3}b_{4}}$ gives cause for concern, as jet vetoes are widely used in experimental analyses at the LHC. For example, vetoing the additional jet activity is crucial for suppressing various SM background processes or in various searches for new physics effects. Indeed, in many analyses carried out by the ATLAS and CMS collaborations the most common jet veto scheme comprises imposing a maximum transverse momentum cut on anti-$k_{T}$ jets. Therefore, in such cases, in order to reduce theoretical uncertainties and to improve the convergence of perturbative predictions, approaches similar to those described in e.g. Ref. Berger:2010xi ; Stewart:2010tn ; Becher:2012qa ; Banfi:2012yh ; Banfi:2012jm ; Michel:2018hui ; Campbell:2023cha would have to be used.
@@
Figure 12: Differential cross-section distributions for the $pp\to t\bar{t}t\bar{t}$ process in the $4\ell$ channel at the LHC with $\sqrt{s}=13.6$ TeV. The upper panels show the absolute NLO QCD predictions for $d\sigma^{\rm NLO}_{\rm exp}/dX$ (blue curve), $d\sigma^{\rm NLO}_{\rm full}/dX$ (red curve) and $d\sigma^{\rm NLO}_{\rm LO_{dec}}/dX$ (green curve), where $X=\Delta\phi_{\ell_{1}\ell_{2}}$ and $\Delta\phi_{b_{1}b_{2}}$. Results are given for $\mu_{R}=\mu_{F}=\mu_{0}=E_{T}/4$ and the NLO MSHT20 PDF set. Also provided are theoretical uncertainties as obtained from the scale dependence. The two middle panels display the ratios of $d\sigma^{\rm NLO}_{\rm full}$ and $d\sigma^{\rm NLO}_{\rm LO_{dec}}$ to $d\sigma^{\rm NLO}_{\rm exp}$ together with their relative NLO scale uncertainties. The lowest panels compare the relative size of NLO scale uncertainties for the three approaches, normalised to their corresponding NLO results. Monte Carlo integration errors are displayed in all panels except the last one.
@@
In the end, we examine the previously defined three approaches $d\sigma^{\rm NLO}_{\rm exp}/dX$, $d\sigma^{\rm NLO}_{\rm full}/dX$ and $d\sigma^{\rm NLO}_{\rm LO_{dec}}/dX$ as a function of the observable $X$, where $X=M_{b_{1}b_{2}},\,p_{T,\,\ell_{1}},\,p_{T,\,b_{1}}$ and $p_{T,\,b_{1}b_{2}}$. These results are shown in Figure 11. The upper panels show the absolute NLO QCD predictions for the three approaches together with their corresponding theoretical uncertainties as obtained from the scale dependence. The two middle panels display the ratios of $d\sigma^{\rm NLO}_{\rm full}/dX$ and $d\sigma^{\rm NLO}_{\rm LO_{dec}}/dX$ to the $d\sigma^{\rm NLO}_{\rm exp}/dX$ result, together with their relative uncertainties. The lowest panels compare again the relative size of NLO scale uncertainties for the three approaches, this time, however, the findings are normalised to their corresponding NLO results for $\mu_{0}=E_{T}/4$. For the three observables studied $M_{b_{1}b_{2}},\,p_{T,\,\ell_{1}}$ and $p_{T,\,b_{1}}$, our observations and conclusions are very similar. The differences between the full results and the expanded ones are of the order of $10\%$, which should be compared to the corresponding NLO uncertainties that are in the range of $15\%-25\%$. On the other hand, NLO QCD corrections in top-quark decays are consistently at the level of $10\%$. Larger effects could be observed for $p_{T,\,b_{1}b_{2}}$, especially in the tail of the distribution. In this case the differences between $d\sigma^{\rm NLO}_{\rm exp}/dp_{T,\,b_{1}b_{2}}$ and $d\sigma^{\rm NLO}_{\rm full}/dp_{T,\,b_{1}b_{2}}$ increase to $20\%$. However, the NLO uncertainties are also increasing and are now in the range of $25\%-35\%$. Nevertheless, the size of higher-order QCD effects in top-quark decays remains similar to the case of the other three observables.
In Figure 12 we additionally display differential cross-section distributions as a function of the difference in azimuthal angle between the two hardest (charged) leptons, $\Delta\phi_{\ell_{1}\ell_{2}}$, and the two hardest $b$-jets, $\Delta\phi_{b_{1}b_{2}}$. The $\Delta\phi_{\ell_{1}\ell_{2}}$ observable is sensitive to signals of numerous beyond the SM scenarios, where among others new heavy states $(Y)$ might be produced in the $pp\to t\bar{t}Y$ process with $Y\to t\bar{t}$ decays. Indeed, angular distributions of charged leptons reflect spin correlations and can be employed to probe the ${\cal CP}$ numbers of $Y$. For $X=\Delta\phi_{\ell_{1}\ell_{2}}$ and $X=\Delta\phi_{b_{1}b_{2}}$ it is interesting to see the difference between $d\sigma^{\rm NLO}_{\rm exp}/dX$ and $d\sigma^{\rm NLO}_{\rm LO_{dec}}/dX$. In both cases we can observe differences in the range of $8\%-11\%$. Furthermore, the differences between the full results and the expanded ones are up to $15\%$, still well within the NLO uncertainties that are of the order of $25\%$. Again, the smallest NLO uncertainties can be seen for the $d\sigma^{\rm NLO}_{\rm full}/dX$ result.
In short, the overall picture is rather similar to that already observed at the integrated fiducial cross-section level. Firstly, the magnitude of the NLO QCD uncertainties increases when higher-order QCD corrections to top-quark decays are neglected. The latter corrections are consistently at the $10\%$ level. Secondly, the differences between the expanded and unexpanded NLO results are within the obtained NLO uncertainties. However, the lack of the expansion of the $1/\Gamma_{t}$ factor in $d\sigma^{\rm NLO}_{\rm full}/dX$ significantly reduces the size of the NLO uncertainties for all the observables that we have examined. Thus, the best NLO QCD predictions for the $pp\to t\bar{t}t\bar{t}$ process in the $4\ell$ channel can be obtained by utilising the $d\sigma^{\rm NLO}_{\rm full}/dX$ results.
7 Summary and Outlook
@@
In the next step, we have compared the following NLO predictions: $d\sigma^{\rm NLO}_{\rm exp}/dX$, $d\sigma^{\rm NLO}_{\rm full}/dX$ and $d\sigma^{\rm NLO}_{\rm LO_{dec}}/dX$ as a function of various observables $X$. By analysing these predictions we have concluded that the overall pattern is rather similar to that which has already been observed at the integrated fiducial cross-section level. Namely, the magnitude of the NLO uncertainties increases when NLO QCD corrections to top-quark decays are neglected. The higher-order corrections in top-quark decays are rather stable and of the order of $10\%$. In addition, the difference between the expanded and unexpanded NLO results is within the obtained NLO scale uncertainties. Nonetheless, the lack of the expansion of the $1/\Gamma_{t}$ term in $d\sigma^{\rm NLO}_{\rm full}/dX$, significantly reduces the size of the NLO uncertainties for all the observables that we have scrutinised. Consequently, the best NLO QCD predictions for the process at hand can be obtained by employing $\sigma^{\rm NLO}_{\rm full}$ and $d\sigma^{\rm NLO}_{\rm full}/dX$.
@@
In addition, it would be advantageous to examine the so-called Matrix Element Corrections (MEC) to the Pythia8 parton shower program Norrbin:2000uu that can upgrade the accuracy of top-quarks decays to approximately NLO QCD in the shower evolution. Quite recently, a study has been carried out taking into account the MEC for the simplest case of $pp\to t\bar{t}$ production in di-lepton top-quark decay channel Frixione:2023hwz . Another study has just been performed for the $pp\to t\bar{t}W^{\pm}$ process for the two same-sign leptons and jets $(2SS\ell)$ as well as three-lepton $(3\ell)$ final states Frederix:2024psm . Unfortunately, such analyses are not yet available for the $pp\to t\bar{t}t\bar{t}+X$ process. Although NLO plus parton shower predictions in the Powheg Box and Mc@@@@Nlo framework for $pp\to t\bar{t}t\bar{t}+X$ production in the $\ell+jets$ channel are available in the literature Jezo:2021smh , the MEC to the top-quark decays are not included there. As this is a much more complicated process with two $t\bar{t}$ pairs and at least 12 final state particles, it would be interesting to test the validity of such approximate methods in such a complex environment. Another very important issue to look at when using the MEC is spin correlations. Having fixed-order predictions that describe top-quark decays and spin correlations with NLO QCD accuracy, we are in an excellent position to test both aspects. Again, we plan to perform such a comparison in the near future.
@@
The obtained NLO QCD results for the $pp\to t\bar{t}t\bar{t}+X$ process are especially important for the High-Luminosity phase of the LHC (HL-LHC). The upgrade is currently in progress and ATLAS and CMS experiments are expected to start taking data from the beginning of 2029. The HL-LHC will allow us to study $pp\to t\bar{t}t\bar{t}$ production in the $4\ell$ channel in greater detail. We point out that our conclusions regarding the magnitude of NLO QCD corrections, the significance of higher-order effects in top-quark decays or the size of NLO uncertainties will remain the same if the center-of-mass energy of the LHC is increased from the current value to $14$ TeV. Finally, we note that, despite its relatively small cross section, a good theoretical control over the $pp\to t\bar{t}t\bar{t}$ process in the $4\ell$ channel is phenomenologically very relevant. Indeed, the NLO QCD uncertainties that we have estimated for this channel are at the level of $15\%-20\%$. For comparison, the projected uncertainties for the HL-LHC for the $4\ell$ channel with ±$\ell^{\pm}=e^{\pm},\mu^{\pm}$, combining ATLAS and CMS, are in the range of $\delta=1/\sqrt{N}=18\%-16\%$. On the other hand, if ±$\ell^{\pm}=e^{\pm},\mu^{\pm},\tau^{\pm}$ the later uncertainties decrease to $8\%-7\%$. We note that these uncertainties are estimated solely on the basis of an integrated luminosity planned for the HL-LHC that amounts to $(3-4)$ ${\rm ab}^{-1}$, see e.g. Ref. ZurbanoFernandez:2020cco . Thus, at the HL-LHC, among other things, we will gain sensitivity to higher-order effects in top-quark decays in the $pp\to t\bar{t}t\bar{t}+X$ process. However, even now, pending new results from HL-LHC, the NLO QCD results obtained in this paper are highly relevant for various preparatory studies. Proper modelling of differential cross-section distributions, for example, is an invaluable contribution to the many Machine Learning studies that are currently being conducted for this process. We also believe that the NLO QCD analysis for the SM $pp\to t\bar{t}t\bar{t}+X$ (background) process at the LHC is a necessary step towards a correct interpretation of possible signals of new physics that may arise in the $4\ell$ channel.
Finally, although the focus of this paper has been on providing NLO QCD results for the $pp\to t\bar{t}t\bar{t}$ process in the $4\ell$ channel in the NWA, and the calculation of higher-order corrections to the full $2\to 12$ process is a huge task at the moment, we can still estimate the impact of non-resonant and off-shell effects at the integrated and differential fiducial cross-section level with the help of the LO predictions. In fact, at the integrated cross-section level the impact of these omitted contributions is of the order of $1\%$ only, and therefore well within the NWA uncertainty defined by ${\cal O}(\Gamma_{t}/m_{t})\approx 0.9\%$, see e.g. Ref. Fadin:1993kt . On the other hand, at the differential cross-section level, neglecting triple-, double-, single- and non-resonant top-quark contributions as well as Breit-Wigner top-quark propagators has a significant impact. In general, full off-shell effects are noticeable for dimensionful observables in the following phase-space regions: in high $p_{T}$ tails and in the vicinity of kinematical endpoints. Since the first type of observables is beyond the reach of the LHC, only the second type should be thoroughly investigated. Indeed, we have checked that differential cross-sectional distributions as a function of the minimum invariant mass of the $b$-jet and the positively charged lepton, $M(\ell^{+}b)$, the reconstructed invariant mass of the top quark, $M(t)$ as well as the stransverse mass of the top quark and $W$ gauge boson, $M_{T2}(t)$ and $M_{T2}(W)$, respectively, are among those observables that are very sensitive to the full off-shell effects in the $pp\to t\bar{t}t\bar{t}$ process. We refer the reader to Ref. Bevilacqua:2022twl for detailed definitions of these observables. In such cases, it would be beneficial to merge the NLO QCD predictions as calculated in the NWA with the LO full off-shell results. To avoid double counting, the quadruple-resonant top-quark contributions must be properly subtracted from the full off-shell result. Such studies are left for the future.